How do I find the average vertical velocity (m/s) of a projectile shooting a ball at 15m/s with a 60 degree angle. I am trying to find its average vertical velocity at 0.4s and 4.41m vertical position.
1
Expert's answer
2020-10-27T11:15:20-0400
"v_y=v\sin{60}-gt\\v_y(0.4s)=15\sin{60}-(9.8)(0.4)=9.07\frac{m}{s}"

"h=vt\sin{60}-0.5gt^2\\4.41=15\sin{60}t-0.5(9.8)t^2\\t=0.4\ s"